# LLM App Store Analysis: A Vision and Roadmap

### v1

@@ 3.1. Security risks
LLM app raw data-related risks. App cloning, where someone unauthorized copies a legitimate app, infringes on intellectual property rights, and potentially introduces security threats or subpar user experiences. In the mobile app ecosystem, app cloning has been a persistent issue, and app stores have employed techniques like code signing (Cooper et al., 2018; Cho et al., 2012) and similarity analysis (Al-Subaihin et al., 2019; Gonzalez et al., 2015) to detect and prevent cloned apps. For LLM app stores, researchers should explore adapting these existing methods, such as code fingerprinting and advanced similarity analysis techniques tailored to LLM apps, to combat app cloning effectively.
@@
App vulnerabilities refer to security weaknesses within LLM apps that attackers can exploit, potentially leading to data breaches or unauthorized activities. For example, these vulnerabilities could arise from inadequate input validation, allowing attackers to perform injection attacks with crafted inputs to elicit unintended responses. This can manipulate LLMs into generating sensitive information, violating content policies, or executing unauthorized actions, compromising app integrity and user data. Additionally, insufficient input checks may render apps vulnerable to jailbreaking, enabling LLMs to output content or perform tasks against terms of service or regulations, raising legal and ethical concerns. These security flaws are also frequently the result of substandard development practices, such as insecure data storage, where sensitive information is poorly protected, making it accessible to unauthorized parties. Weak encryption methods or a lack of robust database security can further exacerbate these issues. Moreover, inadequate authentication mechanisms, including predictable passwords or the absence of multi-factor authentication, can simplify unauthorized access to app functionalities and data. App vulnerabilities are not uncommon in the mobile app ecosystem (Mutchler et al., 2015; Watanabe et al., 2017), with a wealth of established detection techniques available (Abdullah and Zeebaree, 2021; Chen et al., 2022; Shezan et al., 2017). Accordingly, one of the future research directions should be developing tailored solutions to identify and mitigate vulnerabilities specific to LLM apps. This may involve techniques for securing input validation, preventing jailbreaking, enforcing robust authentication, and ensuring secure data storage and transmission within the LLM app ecosystem.
@@
Malicious apps pose a substantial risk in LLM app stores. The malice can manifest in several ways. For example, developers may create LLM apps using instructions or knowledge files containing malicious content, resulting in the app’s knowledge base being tainted with harmful information. Moreover, LLM apps themselves may output content lacking proper constraints, including pornographic or gambling-related information, or even links directing users to malicious websites. Another example is low description-to-behavior fidelity, that is when the actual performance or actions of an app diverge significantly from its documented descriptions or expected behaviors. These phenomena are also prevalent in the mobile app ecosystem. Various tools and techniques have been developed to detect and mitigate malicious mobile apps, such as static and dynamic analysis (Martinelli et al., 2016; Quan et al., 2014; Seo et al., 2014), machine learning-based malware detection (Li et al., 2018; Soni et al., 2020; Wei et al., 2017), and app vetting processes (Backes et al., 2014; Zhang et al., 2013). The unique challenges posed by malicious LLM apps necessitate the development of tailored detection and mitigation strategies. Researchers should focus on developing novel techniques specifically designed to identify and address the distinct threats posed by malicious LLM apps, ensuring a safe and trustworthy ecosystem.
@@
Third-party service integration is another area of concern, as integrating external services or APIs into an app can introduce vulnerabilities or data privacy issues. For example, if the third-party service provider experiences a data breach or has weak security measures, it could compromise the security and privacy of the LLM app and its users. In the mobile app domain, various mature methods have been proposed to address similar issues, including extensive research on third-party library analysis (Feal et al., 2021; Utz et al., 2022; He et al., 2020; Kawabata et al., 2013). To effectively mitigate the risks associated with third-party service integration in LLM apps, developers should adhere to the principle of least privilege, granting only the minimum necessary permissions and access required for the service to function within the app. Robust authentication and authorization mechanisms should be implemented to ensure that only authorized users and processes can interact with the integrated services. Furthermore, encrypting sensitive data both in transit and at rest is crucial when exchanging information with third-party services to protect the confidentiality and integrity of the data. Regular monitoring and auditing of third-party services should also be conducted to detect any suspicious activities or changes in their security posture.
@@
User tracking and profiling without proper consent is another risk, where excessive tracking of user data, behavior, or activities occurs, often for targeted advertising or analyzing user preferences. This can manifest in various harmful ways, such as identity theft, personalized phishing attacks, or unwanted exposure to tailored yet intrusive advertising (Li et al., 2023; Wei et al., 2012; Yalcin, 2017). Moreover, the accumulation and analysis of such data could result in biased or discriminatory outcomes, where decisions made by these LLM apps might favor or disfavor individuals based on their profiled characteristics. This not only undermines user trust but also raises ethical concerns about the fairness and transparency of LLM-powered apps. To mitigate risks associated with user tracking and profiling, LLM app stores should enforce strict privacy policies, obtain explicit user consent, and employ privacy-preserving techniques like differential privacy and data anonymization. Strategies such as regular audits should be adopted to ensure fairness, accountability, and transparency in LLM app decision-making.
@@
Advertisement fraud can occur during the user’s interaction with the LLM app, involving deceptive or misleading ad practices, such as hidden payments, unauthorized data collection, or intrusive ad experiences. Mobile app stores have employed ad network monitoring (Haddadi et al., 2010), real-time ad analysis (Shao et al., 2018), and user feedback analysis (Gui et al., 2015) to combat advertisement fraud. For LLM app stores, researchers should explore adapting these techniques and developing new methods tailored to the unique challenges of LLM apps, ensuring a transparent and trustworthy advertising ecosystem.
@@
Market policy violations, where LLM apps breach the LLM app store’s terms of service, content policies, or other regulations governing app publication and monetization, can undermine the LLM app store’s integrity and user trust. Mobile app stores have implemented automated policy compliance checks (Liu et al., 2021; Zimmeck et al., 2016) and app vetting processes (Zhang et al., 2013) to enforce market regulations. In the context of LLM app stores, researchers should focus on developing automated policy compliance checks tailored to the unique characteristics and challenges of LLM apps, ensuring a secure and trustworthy LLM app ecosystem.
@@
LLM app metadata-related risks. Fake apps, designed to impersonate legitimate LLM apps and deceive users or steal sensitive information, pose a significant risk to users. Mobile app stores have implemented app vetting processes (Zhang et al., 2013) and leveraged techniques like app analysis (Rani et al., 2023) and user feedback (Martens and Maalej, 2019) monitoring to identify fake apps. In the context of LLM app stores, researchers should investigate developing advanced natural language processing and multimedia analysis methods to aid in the detection of fake LLM apps, ensuring user safety and trust.
@@
User feedback-related risks. In the context of LLM app stores, security risks should be carefully considered and addressed. One significant risk is ranking fraud, where attackers attempt to manipulate the LLM app store rankings through illegal methods, such as using bot programs to generate fake ratings, downloads, or reviews, or engaging in keyword stuffing. Similar to the mobile app market, where researchers have proposed systems to detect ranking fraud by analyzing leading sessions, rating patterns, and review behaviors (Zhu et al., 2013, 2014), app stores may need to employ advanced techniques to identify and mitigate fraudulent activities aimed at artificially inflating app rankings and popularity.
@@
Spam reviews in LLM app stores can also contain malicious content or involve large-scale fake reviews manipulated by bots or manual efforts, intending to inflate the app’s reputation artificially. This issue is well-documented in the mobile app industry, where spam reviews and review fraud have been a persistent challenge. Various detection methods have been established in the mobile app domain (Genc, 2019; Seneviratne et al., 2015). Similarly, LLM app stores must adopt and refine such techniques to preserve the authenticity and reliability of their review mechanisms.
@@ 3.2. Privacy protection
Protecting privacy is a critical aspect of LLM app stores. For developers, it is essential to filter out any privacy data that may be included in the instructions or knowledge base files provided to the LLM app. This includes not only personal identifiable information (PII) (Schwartz and Solove, 2011) such as addresses, contact details, and other sensitive data that could compromise individual user privacy, but also extends to sensitive information related to businesses, governmental bodies, and other entities. Protecting this wider range of data ensures the privacy and security of related stakeholders, safeguarding against potential misuse, data breaches, or other forms of exploitation that could have far-reaching consequences. This is similar to the principles and practices adopted in the mobile app industry, where developers are required to implement appropriate data protection measures to safeguard user privacy and comply with relevant regulations, such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017).
Furthermore, developers must also comply with the LLM app store’s privacy policies and relevant legal regulations when collecting user information for personalized fine-tuning or optimization of the app. This involves clearly informing users about the purpose, scope, and manner of data collection and usage, and obtaining user consent. The LLM app store should review the app to ensure compliance with these requirements. Again, this aligns with the standard practices in mobile app stores, where apps are vetted for their data collection and privacy practices, and users are provided with clear information about how their data is being used (Almuhimedi et al., 2015; Shklovski et al., 2014).
From the user’s perspective, privacy data filtering is crucial when interacting with LLM apps. Users’ input may contain private information, and the app should have filtering mechanisms in place to identify and remove this sensitive data, preventing it from being leaked to developers or stored in the knowledge base. This is analogous to the privacy protection measures implemented in mobile apps, where user inputs and data are often processed locally on the device or through secure channels to protect user privacy (Liu et al., 2014; Nema et al., 2022). Additionally, users can expect LLM app stores to provide transparent information about the privacy practices of listed apps, similar to how mobile app stores provide privacy labels and summaries to help users make informed decisions (Apple, 2024b).
@@ 4.1. Developer engagement
Providing comprehensive development assistance, including documentation, examples, and best practices, can lower entry barriers and guide developers in creating high-quality LLM apps. Drawing inspiration from the mobile app domain’s extensive resources, such as Apple’s Human Interface Guidelines (Apple, 2021b) and Google’s Material Design (Google, 2021b), the LLM app ecosystem should create similar guides tailored to conversational AI, covering topics like prompt engineering (Arvidsson and Axell, 2023), context management (Ge et al., 2023), and multi-turn dialogue handling (Wu et al., 2023).
Offering robust analysis and testing tools or frameworks can assist developers in evaluating app performance, identifying vulnerabilities, and optimizing the user experience, ensuring high-quality output. Mobile app development has benefited from tools like Appium (Verma, 2017) and Espresso (Google, 2021a), which have revolutionized automated testing, enabling developers to catch bugs early and ensure app stability. Similarly, the LLM app ecosystem needs to invest in developing testing frameworks that can simulate user interactions and detect potential biases or inconsistencies in the generated output.
@@ 4.2. Competitive landscape
Leveraging user preferences, history, and ratings, LLM app stores can develop sophisticated recommendation algorithms to suggest potentially interesting LLM apps to users. This not only improves app discoverability but also increases user satisfaction and engagement. Mobile app stores have successfully implemented such recommendation systems, with examples like Apple’s app store featuring “Apps You Might Like” and Google Play’s “Recommended for You” sections (Tu et al., 2019; Hwang et al., 2019). However, in the context of LLM app stores, academic research on tailoring recommendation algorithms to this novel domain remains largely unexplored.
To help developers effectively promote their LLM apps, LLM app stores can offer a range of promotion tools and channels. Mobile app stores usually use ads, featured spots, and events to promote LLM apps, leveraging influencer partnerships for added trust and visibility (Lee et al., 2018; Rahman et al., 2019). Similarly, LLM app stores could also include advertising placement options, such as sponsored search results or featured app listings, allowing developers to increase their app’s visibility to potential users. Additionally, LLM app stores can provide promotional opportunities through curated collections, themed showcases, or developer spotlights, highlighting noteworthy LLM apps and their creators.
Curating and featuring top LLM apps is a strategic move by LLM app stores to influence the app market landscape (Hunter, 2024; OpenAI, 2024a; GPTStore.AI, 2024). Through spotlighting apps that excel in innovation and quality, LLM app stores could establish benchmarks and inspire developers to aim high. This not only guides users to superior apps but also rewards developers for outstanding user experiences.
To improve LLM app discoverability, LLM app stores are encouraged to deploy a coherent app classification system. Sorting LLM apps by their functionality, usage scenarios, industries, or target audiences simplifies the search process for users. This not only elevates the user experience but also supports developers in strategically showcasing their apps. For instance, renowned stores like Google Play (Fu et al., 2013; Google, 2024), Apple App Store (Apple, 2024a; Jia et al., 2020), and Blackberry World App Store (Harman et al., 2012) typically employ a consolidated category structure.

### v2

@@ 4.1. Security risks
LLM app raw data-related risks. App cloning, where someone unauthorized copies a legitimate app, infringes on intellectual property rights and potentially introduces security threats or subpar user experiences. In the mobile app ecosystem, app cloning has been a persistent issue, and app stores have employed techniques like code signing (Cooper et al., 2018; Cho et al., 2012) and similarity analysis (Al-Subaihin et al., 2019; Gonzalez et al., 2014) to detect and prevent cloned apps. However, in the context of LLM apps, cloning challenges differ due to the reliance on proprietary base LLMs and unique prompt engineering strategies. This necessitates the development of specialized detection mechanisms that account for these distinctive features, such as monitoring for duplicated prompt patterns or misuse of LLMs. For LLM app stores, researchers should explore similarity analysis techniques tailored to LLM apps, to combat app cloning effectively.
@@
App vulnerabilities refer to security weaknesses within LLM apps that attackers can exploit, potentially leading to data breaches or unauthorized activities. For example, these vulnerabilities could arise from inadequate input validation, allowing attackers to perform injection attacks with crafted inputs to elicit unintended responses. Unlike traditional mobile apps, LLM apps are particularly susceptible to prompt injection attacks and adversarial inputs that exploit the language understanding capabilities of LLMs, requiring novel defense mechanisms specifically designed to address these challenges. This can manipulate LLMs into generating sensitive information, violating content policies, or executing unauthorized actions, compromising app integrity and user data. Additionally, insufficient input checks may render apps vulnerable to jailbreaking, enabling LLMs to output content or perform tasks against terms of service or regulations, raising legal and ethical concerns.
These security flaws are also frequently the result of substandard development practices, such as insecure data storage, where sensitive information is poorly protected, making it accessible to unauthorized parties. Weak encryption methods or a lack of robust database security can further exacerbate these issues. Moreover, inadequate authentication mechanisms, including predictable passwords or the absence of multi-factor authentication, can simplify unauthorized access to app functionalities and data. App vulnerabilities are not uncommon in the mobile app ecosystem (Mutchler et al., 2015; Watanabe et al., 2017), with a wealth of established detection techniques available (Abdullah and Zeebaree, 2021; Chen et al., 2022; Shezan et al., 2017). Accordingly, one of the future research directions should be developing tailored solutions to identify and mitigate vulnerabilities specific to LLM apps. This may involve techniques for securing input validation, preventing jailbreaking, enforcing robust authentication, and ensuring secure data storage and transmission within the LLM app ecosystem.
@@
Malicious apps pose a substantial risk in LLM app stores. The malice can manifest in several ways. For example, developers may create LLM apps using instructions or knowledge files containing malicious content, resulting in the app’s knowledge base being tainted with harmful information. Moreover, LLM apps themselves may output content lacking proper constraints, including pornographic or gambling-related information, or even links directing users to malicious websites. Another example is low description-to-behavior fidelity, that is when the actual performance or actions of an app diverge significantly from its documented descriptions or expected behaviors. These phenomena are also prevalent in the mobile app ecosystem. Various tools and techniques have been developed to detect and mitigate malicious mobile apps, such as static and dynamic analysis (Martinelli et al., 2016; Quan et al., 2014; Seo et al., 2014), machine learning-based malware detection (Li et al., 2018; Soni et al., 2020; Wei et al., 2017), and app vetting processes (Backes et al., 2014; Zhang et al., 2013). Given that LLM apps can dynamically generate content, malicious actors might exploit this capability to produce harmful outputs that are harder to detect using traditional static analysis, emphasizing the need for advanced real-time monitoring and content validation methods unique to LLM environments. The unique challenges posed by malicious LLM apps necessitate the development of tailored detection and mitigation strategies. Researchers should focus on developing novel techniques specifically designed to identify and address the distinct threats posed by malicious LLM apps, ensuring a safe and trustworthy ecosystem.
@@
Third-party service integration is another area of concern, as integrating external services or APIs into an app can introduce vulnerabilities or data privacy issues. For example, if the third-party service provider experiences a data breach or has weak security measures, it could compromise the security and privacy of the LLM app and its users. In the mobile app domain, various mature methods have been proposed to address similar issues, including extensive research on third-party library analysis (Feal et al., 2021; Utz et al., 2023; He et al., 2022; Kawabata et al., 2013). To effectively mitigate the risks associated with third-party service integration in LLM apps, developers should adhere to the principle of least privilege, granting only the minimum necessary permissions and access required for the service to function within the app. Robust authentication and authorization mechanisms should be implemented to ensure that only authorized users and processes can interact with the integrated services. Furthermore, encrypting sensitive data both in transit and at rest is crucial when exchanging information with third-party services to protect the confidentiality and integrity of the data. Regular monitoring and auditing of third-party services should also be conducted to detect any suspicious activities or changes in their security posture. Additionally, the interactive and adaptive nature of LLM apps may inadvertently expose sensitive user interactions to third-party services, necessitating more stringent data handling policies and privacy-preserving integrations specifically designed for the LLM context.
@@
User tracking and profiling without proper consent is another risk, where excessive tracking of user data, behavior, or activities occurs, often for targeted advertising or analyzing user preferences. This can manifest in various harmful ways, such as identity theft, personalized phishing attacks, or unwanted exposure to tailored yet intrusive advertising (Li et al., 2023; Wei et al., 2012; Yalcin, 2017). Unlike conventional apps, LLM apps may infer additional personal information from seemingly innocuous user inputs due to their advanced language understanding, amplifying privacy concerns and highlighting the need for specialized safeguards against unintended data collection and profiling. Moreover, the accumulation and analysis of such data could result in biased or discriminatory outcomes, where decisions made by these LLM apps might favor or disfavor individuals based on their profiled characteristics. This not only undermines user trust but also raises ethical concerns about the fairness and transparency of LLM-powered apps. To mitigate risks associated with user tracking and profiling, LLM app stores should enforce strict privacy policies, obtain explicit user consent, and employ privacy-preserving techniques like differential privacy and data anonymization. Strategies such as regular audits should be adopted to ensure fairness, accountability, and transparency in LLM app decision-making.
@@
Advertisement fraud can occur during the user’s interaction with the LLM app, involving deceptive or misleading ad practices, such as hidden payments, unauthorized data collection, or intrusive ad experiences. Mobile app stores have employed ad network monitoring (Haddadi et al., 2010), real-time ad analysis (Shao et al., 2018), and user feedback analysis (Gui et al., 2015) to combat advertisement fraud. However, LLM apps can dynamically generate content, including advertisements, which may be manipulated to display fraudulent or malicious ads that are harder to detect using conventional monitoring techniques. This necessitates advanced AI-driven solutions that can analyze and interpret generated content in real-time, specifically tailored to the generative capabilities of LLMs. For LLM app stores, researchers should explore adapting these techniques and developing new methods tailored to the unique challenges of LLM apps, ensuring a transparent and trustworthy advertising ecosystem.
@@
Market policy violations, where LLM apps breach the LLM app store’s terms of service, content policies, or other regulations governing app publication and monetization, can undermine the LLM app store’s integrity and user trust. Mobile app stores have implemented automated policy compliance checks (Liu et al., 2021; Zimmeck et al., 2017) and app vetting processes (Zhang et al., 2013) to enforce market regulations. Due to the generative and adaptive capabilities of LLMs, policy violations may occur dynamically during user interactions, requiring continuous monitoring and compliance enforcement mechanisms that are specifically designed for the evolving outputs of LLM apps. In the context of LLM app stores, researchers should focus on developing automated policy compliance checks tailored to the unique characteristics and challenges of LLM apps, ensuring a secure and trustworthy LLM app ecosystem.
@@
LLM app metadata-related risks. Fake apps, designed to impersonate legitimate LLM apps and deceive users or steal sensitive information, pose a significant risk to users. Mobile app stores have implemented app vetting processes (Zhang et al., 2013) and leveraged techniques like app analysis (Rani et al., 2023) and user feedback (Martens and Maalej, 2019) monitoring to identify fake apps. Given the sophisticated language capabilities of LLMs, fake apps can produce highly convincing content and interactions, making them more deceptive than traditional fake apps and requiring more nuanced detection strategies that consider the quality and context of generated outputs. In the context of LLM app stores, researchers should investigate developing advanced natural language processing and multimedia analysis methods to aid in the detection of fake LLM apps, ensuring user safety and trust.
@@
User feedback-related risks. In the context of LLM app stores, security risks should be carefully considered and addressed. One significant risk is ranking fraud, where attackers attempt to manipulate the LLM app store rankings through illegal methods, such as using bot programs to generate fake ratings, downloads, or reviews, or engaging in keyword stuffing. Similar to the mobile app market, where researchers have proposed systems to detect ranking fraud by analyzing leading sessions, rating patterns, and review behaviors (Zhu et al., 2013, 2015), app stores may need to employ advanced techniques to identify and mitigate fraudulent activities aimed at artificially inflating app rankings and popularity. Moreover, attackers might leverage LLMs themselves to generate human-like fake reviews and feedback at scale, complicating detection efforts and necessitating the development of sophisticated algorithms capable of distinguishing between genuine and AI-generated content.
@@
Spam reviews in LLM app stores can also contain malicious content or involve large-scale fake reviews manipulated by bots or manual efforts, intending to inflate the app’s reputation artificially. This issue is well-documented in the mobile app industry, where spam reviews and review fraud have been a persistent challenge. Various detection methods have been established in the mobile app domain (Genc, 2019; Seneviratne et al., 2015). Similarly, LLM app stores must adopt and refine such techniques to preserve the authenticity and reliability of their review mechanisms. Considering that LLMs can generate high-quality text, spam reviews may become more sophisticated to distinguish from genuine user feedback, requiring enhanced detection methods that analyze linguistic patterns and contextual relevance specific to LLM-generated content.
@@ 4.2. Privacy protection
Protecting privacy is a critical aspect of LLM app stores. For developers, it is essential to filter out any privacy data that may be included in the instructions or knowledge base files provided to the LLM app. This includes not only personally identifiable information (PII) (Schwartz and Solove, 2011) such as addresses, contact details, and other sensitive data that could compromise individual user privacy, but also extends to sensitive information related to businesses, governmental bodies, and other entities. Protecting this wider range of data ensures the privacy and security of related stakeholders, safeguarding against potential misuse, data breaches, or other forms of exploitation that could have far-reaching consequences. This is similar to the principles and practices adopted in the mobile app industry, where developers are required to implement appropriate data protection measures to safeguard user privacy and comply with relevant regulations, such as the General Data Protection Regulation (GDPR) (Voigt and Von dem Bussche, 2017).
Furthermore, developers must also comply with the LLM app store’s privacy policies and relevant legal regulations when collecting user information for personalized fine-tuning or optimization of the app. This involves clearly informing users about the purpose, scope, and manner of data collection and usage, and obtaining user consent. The LLM app store should review the app to ensure compliance with these requirements. Again, this aligns with the standard practices in mobile app stores, where apps are vetted for their data collection and privacy practices, and users are provided with clear information about how their data is being used (Almuhimedi et al., 2015; Shklovski et al., 2014). Given that LLM apps may process large volumes of unstructured user-generated text, which can inadvertently contain sensitive information, developers must implement advanced consent mechanisms and data handling policies that address the specific risks associated with natural language data processing, ensuring that users are adequately informed and protected.
From the user’s perspective, privacy data filtering is crucial when interacting with LLM apps. Users’ input may contain private information, and the app should have filtering mechanisms in place to identify and remove this sensitive data, preventing it from being leaked to developers or stored in the knowledge base. This is analogous to the privacy protection measures implemented in mobile apps, where user inputs and data are often processed locally on the device or through secure channels to protect user privacy (Liu et al., 2014; Nema et al., 2022). However, the conversational and predictive capabilities of LLMs mean that they might generate outputs based on sensitive user data, even when not explicitly provided in a single interaction, requiring innovative privacy safeguards that prevent the unintentional disclosure of personal information through model predictions. Additionally, users can expect LLM app stores to provide transparent information about the privacy practices of listed apps, similar to how mobile app stores provide privacy labels and summaries to help users make informed decisions (Apple, 2024b).
@@ 5.1. Developer engagement
Providing comprehensive development assistance, including documentation, examples, and best practices, can lower entry barriers and guide developers in creating high-quality LLM apps. Drawing inspiration from the mobile app domain’s extensive resources, such as Apple’s Human Interface Guidelines (Apple, 2021b) and Google’s Material Design (Google, 2021b), the LLM app ecosystem should create similar guides tailored to conversational AI, covering topics like prompt engineering (Arvidsson and Axell, 2023), context management (Ge et al., 2024), and multi-turn dialogue handling (Wu et al., 2023).
Offering robust analysis and testing tools or frameworks can assist developers in evaluating app performance, identifying vulnerabilities, and optimizing the user experience, ensuring high-quality output. Mobile app development has benefited from tools like Appium (Verma, 2017) and Espresso (Google, 2021a), which have revolutionized automated testing, enabling developers to catch bugs early and ensure app stability. Similarly, the LLM app ecosystem needs to invest in developing comprehensive testing frameworks that can effectively simulate user interactions and detect potential biases or inconsistencies in the generated output, thus improving overall reliability and user satisfaction.
@@ 5.2. Competitive landscape
Leveraging user preferences, history, and ratings, LLM app stores can develop sophisticated recommendation algorithms to suggest potentially interesting LLM apps to users. This not only improves app discoverability but also increases user satisfaction and engagement. Mobile app stores have successfully implemented such recommendation systems, with examples like Apple’s app store featuring “Apps You Might Like” and Google Play’s “Recommended for You” sections (Tu et al., 2020; Hwang et al., 2019). However, in the context of LLM app stores, academic research on tailoring recommendation algorithms to this novel domain remains largely unexplored.
To help developers effectively promote their LLM apps, LLM app stores can offer a range of promotion tools and channels. Mobile app stores usually use ads, featured spots, and events to promote mobile apps, leveraging influencer partnerships for added trust and visibility (Lee et al., 2018; Rahman et al., 2019). Similarly, LLM app stores could also include advertising placement options, such as sponsored search results or featured app listings, allowing developers to increase their app’s visibility to potential users. Additionally, LLM app stores can provide promotional opportunities through curated collections, themed showcases, or developer spotlights, highlighting noteworthy LLM apps and their creators.
Curating and featuring top LLM apps is a strategic move by LLM app stores to influence the app market landscape (Hunter, 2024; OpenAI, 2024; GPTStore.AI, 2024). LLM app stores could establish benchmarks and inspire developers to aim high by spotlighting apps that excel in innovation and quality. This not only guides users to superior apps but also rewards developers for outstanding user experiences.
LLM app stores are encouraged to deploy a coherent app classification system to improve LLM app discoverability. Sorting LLM apps by their functionality, usage scenarios, industries, or target audiences simplifies the search process for users. This not only elevates the user experience but also supports developers in strategically showcasing their apps. For instance, renowned stores like Google Play (Fu et al., 2013; Google, 2024), Apple App Store (Apple, 2024a; Jia et al., 2020), and Blackberry World App Store (Harman et al., 2012) typically employ a consolidated category structure.